OpenStudy (anonymous):

simplify sqrt(48x^2/25)

OpenStudy (anonymous):

4sqrt3*x/5

OpenStudy (anonymous):

(4x/5)*sqrt(3)

OpenStudy (anonymous):

\[\sqrt{\frac{48x^2}{25}}=\sqrt{\frac{16\cdot3x^2}{25}}=\frac{4x}{5}\sqrt{3}\]

OpenStudy (anonymous):

4x53√3 is correct